yeetbeet420:

Which ordered pair is the solution to the system of linear equations Negative 5 x + y = 26 and 2 x minus 7 y = 16? (–4, 6) (6, –4) (–4, –6) (–6, –4)

justus:

y = 5x +26 2x -7(5x +26) = 16 2x -35x -182 = 16 -198 = 33x then add 33x-16 -6 = x divide by 33 y = 5(-6) +26 = -4

yeetbeet420:

I beleve it is (–6, –4) but idk

justus:

Correct!

justus:

:)

yeetbeet420:

nice

yeetbeet420:

thx

justus:

NP!